OpenStudy (anonymous):

Could you please help me with this @ganeshie8 ??? I need help to figure this out :D thx! Based on simple interest, find the total amount saved or owed. principal = $1056; rate = 3.75%; time = 6 years

ganeshie8 (ganeshie8):

so you have deposited $1056 in bank

ganeshie8 (ganeshie8):

and bank gives 3.75% interest on ur money every year

ganeshie8 (ganeshie8):

use below formula for total interest u wud get :- \(\large Total ~ Interest = \frac{P*R*T}{100}\)

ganeshie8 (ganeshie8):

\(\large Total ~ Interest = \frac{1056~*~3.75~*~6}{100}\)

ganeshie8 (ganeshie8):

simplify, wat do u get ?

OpenStudy (anonymous):

@ganeshie8 now what?

ganeshie8 (ganeshie8):

im not sure i understand ur table above :/

ganeshie8 (ganeshie8):

\(\large Total ~ Interest = \frac{1056~*~3.75~*~6}{100}\) \(\large Total ~ Interest = 237.6 \)

ganeshie8 (ganeshie8):

thats the total Interest you wud get in 6 years okay ?

OpenStudy (anonymous):

ok :( I still don't get it D;

ganeshie8 (ganeshie8):

add that interest to principal, total money in bank = 1056 + 237.6 = 1293.6

OpenStudy (anonymous):

how did u get 237.6?

ganeshie8 (ganeshie8):

i have used calculator : http://www.wolframalpha.com/input/?i=%5Cfrac%7B1056*3.75*6%7D%7B100%7D

ganeshie8 (ganeshie8):

:)

OpenStudy (anonymous):

@ganeshie8 that's not a possible answer

ganeshie8 (ganeshie8):

oh u have options ha ?

ganeshie8 (ganeshie8):

we must have done mistake somewhere... check once :)

OpenStudy (anonymous):

yes hehe

ganeshie8 (ganeshie8):

lol wat are the options

OpenStudy (anonymous):

ugh! I lost the probalm bc my internet went out lol...but heres one like it _________________________________________________________________________ Find the simple interest. principal = $10,400; rate = 12.5%; time = 11 years A. $1300 B. $9152 C. $11,700 D. $14,300

OpenStudy (anonymous):

sorry for all the spacing in-between lol...I copied and pasted it d:

ganeshie8 (ganeshie8):

Use the same formula

ganeshie8 (ganeshie8):

\(\large Simple ~interest = \frac{P~*~R~*~T}{100}\)

ganeshie8 (ganeshie8):

\(\large Simple ~interest = \frac{10400~*~12.5~*~11}{100}\)

ganeshie8 (ganeshie8):

simplify, you would get 14,300 : http://www.wolframalpha.com/input/?i=10400*12.5*11+%2F100

OpenStudy (anonymous):

OH!!! I get it now!!!!!!! you times them all then just divide by 100! :D :D :D!! THX!

ganeshie8 (ganeshie8):

yup :)

OpenStudy (anonymous):

...ok it didn't work....I tried it again.. with another problem.. and the answer I got was not a possible anwer.. :(

ganeshie8 (ganeshie8):

whats the question ?

OpenStudy (anonymous):

OH!! its a diff questions.. the first was 1.Find the simple interest. 2.Based on simple interest, find the total amount saved or owed

OpenStudy (anonymous):

Based on simple interest, find the total amount saved or owed. principal = $1056; rate = 3.75%; time = 6 years A. $1293.60 B. $1452.00 C. $1689.00 D. $3432.00

ganeshie8 (ganeshie8):

find the simple interest first

ganeshie8 (ganeshie8):

simple interest = (1056 * 3.75 * 6) / 100 = ?

OpenStudy (anonymous):

=23,760

ganeshie8 (ganeshie8):

Nope

OpenStudy (anonymous):

:(

ganeshie8 (ganeshie8):

you should get "237.6" http://www.wolframalpha.com/input/?i=+%281056+*+3.75+*+6%29+%2F+100+

ganeshie8 (ganeshie8):

thats the simple interest

OpenStudy (anonymous):

=237.6...ohh I was typing when u sent that...I forgot to divide by 100

ganeshie8 (ganeshie8):

:)

ganeshie8 (ganeshie8):

total money saved = 1056 + 237.6 = ?

OpenStudy (anonymous):

=1293.6?

OpenStudy (anonymous):

so..A!

ganeshie8 (ganeshie8):

You got it !

OpenStudy (anonymous):

YAY!

ganeshie8 (ganeshie8):

good job :)

OpenStudy (anonymous):

:D hwo about this one...how do I do it? The simple interest on an investment of $2000 for one year is $110. What is the interest rate? A. 0.0055% B. 0.055% C. 0.55% D. 5.5%

ganeshie8 (ganeshie8):

use te same formula, but this time u need to solve for \(R\)

ganeshie8 (ganeshie8):

I = P*T*T/100

ganeshie8 (ganeshie8):

110 = 2000*1*R/100

ganeshie8 (ganeshie8):

solve R